OpenStudy (jasmineflvs):

Hey guys! My name is Jasmine! This isn't a question, but I just wanted to say if anyone needs help on geometry related things that I would be honored to help. Just shoot me a message or tag me and I'll be willing to help, because I love geometry and I don't see very many people that ask for help on it here.

OpenStudy (bibby):

if a rectangle is a parallelogram and a triangle has 3 sides will you be my girlfriend

OpenStudy (jasmineflvs):

hahahaha that's funny.

OpenStudy (bibby):

it was worth a shot my mom really wants grandchildren you see I'll stop harassing you now rofl

OpenStudy (jasmineflvs):

ahaha!!!! well hopefully you find a geometry loving girl in your world.

OpenStudy (anonymous):

funniest talk ever

OpenStudy (bohotness):

hi i have the same name as you

OpenStudy (demonchild99):

lol....ok my sister is doin the Pythagorean therum...i tried to explain that if she does a^2 +b^2 =c^2 so if a=9 b=12 she doesnt understand that you have to square them...she is trying to times them by two..how can i tell her she has to multiply it by the same number..lol..is there an easier way to explain that?

OpenStudy (jasmineflvs):

oooooooo hmm...well tell her that its like copying the same number to the same number and just keep copying it...but multiplying it by two is just adding that number to its self one time....like 3 times 2 would be 6 ....meaning 3+3 is six but 3 ^2 would be 3 times 3 which would be nine...3+3+3 =9

OpenStudy (cookielate):

hi can you help me make a visual pattern please

OpenStudy (anonymous):

what is the area of each sail? Which sailboat has the greatest total sail area? first sailboat 6ft h and 9 ft width. the second sailboat has two different triangles 5 ft h and 7 ft width, 5 ft h and 3 ft width.

OpenStudy (jasmineflvs):

hmmm I don't know that one for sure

OpenStudy (jasmineflvs):

@bohotness oh really cool

OpenStudy (demonchild99):

@lilycat i would say just lookin at the numbers the first sail has the biggest over all area..i cannot tell you exactly what that area is tho...im sorry

OpenStudy (serenity74):

creepiest talk ever.................#creepedoutforlife

OpenStudy (bibby):

It's not creepy, it's romantic :/

OpenStudy (demonchild99):

wow i cant believe anyone found this haha...this was from a whole year ago o my goodness